OpenStudy (anonymous):

Solve the following system: will medal! 5x - 2y = -11 -6x + 4y = 10

OpenStudy (anonymous):

solution (x,y)= ( , )

OpenStudy (anonymous):

lets do this by subsititution lets find x 2(5x-2y=-11) -6x+4y=10 10x-4y=-22 -6x+4y=10 4x=-32 x=-8

OpenStudy (apolloschariot):

@ilikemath50 , it is against the CoC to just give answers and your account could be suspended for it

OpenStudy (anonymous):

6(5x-2y=-11) 5(-6x+4y=10) 30x-12y=-66 -30x+20=10 8y=-56 y=-7

OpenStudy (anonymous):

no its not everyone does it

OpenStudy (anonymous):

(-8,-7)

OpenStudy (anonymous):

it is against the CoC but ppl ignore it

OpenStudy (anonymous):

thanks what about this

OpenStudy (anonymous):

5x - 2y = -11 7x - 3y = -15

OpenStudy (anonymous):

whats coc?

OpenStudy (anonymous):

code of conduct

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

-7(5x-2y=-11) 5(7x-3y=-15) -35x+14y=77 35x-15y=-75 -y=2 y=-2

OpenStudy (anonymous):

-8,-2?

OpenStudy (anonymous):

and this is elimination not substitution

OpenStudy (anonymous):

-3(5x-2y=-11) 2(7x-3y=-15) -15x+6y=33 14x-6y=-30 -x=3 x=-3

OpenStudy (anonymous):

is that an option? -3,-2

OpenStudy (anonymous):

no i have to just write it no mult choice

OpenStudy (anonymous):

ahh alright it's -3,-2